OpenStudy (anonymous):

John di Silvestre obtained a loan to buy a cello for $8,400. He pays 12% exact interest on the loan for 146 days. What is the amount of interest owed and the maturity value of the loan? A) $1,680.00 interest; $10,080.00 maturity value B) $1,008.00 interest; $9,408.00 maturity value C) $3,360.00 interest; $11,360.00 maturity value D) $403.20 interest; $8,803.20 maturity value

OpenStudy (ranga):

Normally when they tell you interest rate they say per annum meaning the percent interest rate for a year. But in this problem they seem to say John pays 12% for a 146 day loan. So interest = P * r / 100 = 8400 * 12/100 = $1,008.00 Maturity value = Principal + Interest = 8400 + 1008 = $9,408.00